The equation ax^2 + bx + c = 0 can be solved using fromula method and this is a quadratic in x. my question is how to make x as subject in the equation
y= (4/x^2) + (5x + 8/x^2)

Answer by pushpaharan(47)

what do you mean by " how to make x as subject in the equation ".
also question seems vague. is it y= (4/x^2) + (5x + 8/x^2) correct.
y= (4/x^2) + (5x + 8/x^2)
= (12/x^2) + 5x
= (12 + 5 x^3)/(x^2)
Please post the question again if the equation is wrong.
thanks